Shift report 20030113 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Mon Jan 13 2003 - 00:40:50 EST

  • Next message: Ian Bearden: "RE: Shift report 20030113 16:00-24:00"
    ------------------------------------------------------------
    Date: Mon Jan 13 00:40:50 2003, Shift: 16:00-24:00
    Supervisor: ykl, with: cr
    Log book#13 pages 54 - 56
    
    Shift Summary:
    ------------------------------------------------------------
    1. One store in the afternoon with low d intensity and
    on long store in the evening with a high d intensity.
    
    2.Reached the goal for 250K trigger 6 at FS setting of 1/8A
    by run7347. Started  FS setting of 1/4A by run7349.
    
    SJS and JHL worked on trigger timing and eliminated the
    noise peak in vertexz spectrum.
    
    3.T3 is still turned off. T1, DVM, and RICH are all
    working now. 
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Mon Jan 13 2003 - 00:45:23 EST